OpenStudy (mikey):

Find the solution of the square root of the quantity of x plus 2 plus 4 equals 6, and determine if it is an extraneous solution. x = 2; extraneous x = 2; not extraneous x = 30; extraneous x = 30; not extraneous

OpenStudy (mikey):

I know that its x=2 but not sure if its extraneous or not

563blackghost (563blackghost):

Solve the problem again... \(\huge{\sqrt{x+2+4}=6}\) First raise both by power... \(\huge{x+6=36}\) What would x equal?

563blackghost (563blackghost):

@mikey

OpenStudy (mikey):

I thought u would subtract four first

OpenStudy (mikey):

but since you did it that way it would be x=30

563blackghost (563blackghost):

No we would not subtract since this is with a square root....in order to get rid of the square root we would first need to square each side of the equation..... square root by squared is canceled out and we do 6 squared which would give us 36....we also simplify 2 and 4 to be 6... Correct :)

563blackghost (563blackghost):

No extraneous solution are solution given by simplifying but will not equal the total....now simplify the equation with x inputted.... \(\huge{\sqrt{30+2+4}=6}\) Does \(\Large{\sqrt{30+2+4}}\) equal `6`?

563blackghost (563blackghost):

Sorry not `No` i meant `Now`...

OpenStudy (mikey):

yes they are equal

563blackghost (563blackghost):

So `x=30` is not an extraneous solution :)

OpenStudy (mikey):

k thank you soo much

563blackghost (563blackghost):

np ;)